Solution A contains 40% of concentrated acid and SolutionB contains 60% of the same concentrated acid. How many cubic acid centimetres of each type are needed to produce 1000cm^3 of mixture containing 55% concentrated acid?
Answer by robertb(5830)

Let x = # cc of the 40% concentrate.
Then 1000 - x = #cc of the 60 concentrate.
==> 0.40x + 0.60(1000- x) = 0.55(1000)
0.40x + 600 - 0.60x = 550
-x/5 = -50
x = 250 cc of the 40% concentrate, and
750 cc of the 60% concentrate.